# UMArm: Untethered, Modular, Wearable, Soft Pneumatic Arm

### v1

@@ Abstract
Robotic arms are essential to modern industries, however, their adaptability to unstructured environments remains limited. Soft robotic arms, particularly those actuated pneumatically, offer greater adaptability in unstructured environments and enhanced safety for human-robot interaction. However, current pneumatic soft arms are constrained by limited degrees of freedom, precision, payload capacity, and reliance on bulky external pressure regulators. In this work, a novel pneumatically driven rigid-soft hybrid arm, “UMArm”, is presented. The shortcomings of pneumatically actuated soft arms are addressed by densely integrating high-force-to-weight-ratio, self-regulated McKibben actuators onto a lightweight rigid spine structure. The modified McKibben actuators incorporate valves and controllers directly inside, eliminating the need for individual pressure lines and external regulators, significantly reducing system weight and complexity. Full untethered operation, high payload capacity, precision, and directionally tunable compliance are achieved by the UMArm. Portability is demonstrated through a wearable assistive arm experiment, and versatility is showcased by reconfiguring the system into an inchworm robot. The results of this work show that the high-degree-of-freedom, external-regulator-free pneumatically driven arm systems like the UMArm possess great potential for real-world unstructured environments.
Index Terms:
Soft robotic arms, Pneumatic actuators, Untethered systems, Wearable device
@@ I Introduction
Robotic arms are essential components in modern industries. Rigid robotic arms can operate with great precision, speed, and payload capacity, making them ideal for highly structured environments like factories [1, 2, 3]. However, their application in unstructured real-world environments remains limited due to their lack of compliance. In unstructured environments, compliance is crucial for safe interaction and effective handling of unexpected disturbances. Compliance can be achieved either by structure or by control. Rigid robots can simulate compliance through control, but it requires the coordinated control of all actuators, and will be limited by controller bandwidth and sensor latency. Without full-body compliance, rigid arms have to operate on highly precise and low-latency spatial information—a requirement that is difficult to satisfy in unstructured environments [4, 5]. Consequently, rigid arms can cause damage to target objects, nearby humans, or the robot itself [6, 7, 8].
@@
Unlike their rigid counterparts, soft robotic arms have structural compliance. They are capable of mitigating external disturbances and compensating for misalignment through elastic deformation of their structure, making them more suitable for operation in unstructured environments [9, 10, 11]. Their inherent compliance also ensures safety in physical human-robot interaction, enabling use cases such as daily assistance and elder care [12, 13].
@@
A variety of soft actuation strategies have been explored to enable compliance and safety in unstructured environments. HASEL actuators achieve high strain and power output via electrohydraulic zipping, enabling fast, modular actuation; however, their high-voltage requirements and structural fragility pose challenges for unstructured deployment [14]. Shape memory alloy (SMA)-driven soft actuators offer compact and lightweight actuation with high power-to-weight ratios and low voltage requirements but suffer from slow thermal response, nonlinear behavior, and sensitivity to ambient temperature—factors that hinder precise control [15]. Hydraulic soft actuators provide high force output[16], yet their practicality is limited by their bulky fluidic components and slow responses. Tendon-driven designs have also gained attention, offering enhanced workspaces and tunable compliance through mechanisms like trimmed helicoids [17], but these systems remain vulnerable to friction, slack, and lack robustness under dynamic conditions [17, 15].
@@
Among various soft actuation strategies, pneumatic actuators are promising due to their smooth, continuous deformation, high force-to-weight ratio, and compliance in every direction. Unlike cable-driven mechanisms, pneumatic actuators do not require complex cable routing, which simplifies design and thus enhances reliability. Compared to dielectric elastomers, HASEL actuators, and shape memory alloys, pneumatic actuators offer a more practical path toward building human-arm-scale systems that avoids the need for high-voltage operation, complex material packaging, or thermally limited cycle rates. These properties, combined with their resilience to impact, ease of fabrication, and cost-effectiveness, make pneumatic actuators particularly well-suited for building robot arms for deploying into real-world, unstructured environments [18, 19, 20, 21, 22, 23].
@@
Although pneumatically actuated soft arms have desirable characteristics, certain drawbacks limit their performance. Firstly, although many pneumatic soft arms are lightweight, they require bulky external pressure regulators [24]. The cumbersome external hardware limits their portability. Secondly, high dexterity pneumatic soft arms need many independently actuated degrees of freedom (DoF), but the centralized, external regulating scheme makes it difficult to increase independent DoFs. This is because adding independently controlled pneumatic actuators requires extra external regulators and more tubing running from the actuators to the external regulators. Large amounts of tubing will introduce extra weight, thereby affecting the payload capacity of the arm. At the same time, long thin tubes increase fluidic resistance, causing pressure to propagate more slowly from the regulator to the actuator, affecting pressure control performance [25]. Thirdly, structurally soft pneumatic arms might suffer from uncontrolled compliant degrees of freedom (DoF). Under heavy load, they might buckle at unwanted positions, affecting precision and payload capacity [26]. Moreover, robots might oscillate at the uncontrolled compliant joints when given aggressive control inputs [27].
@@
Researchers have explored various pneumatic soft arm architectures to address common limitations. Continuum soft pneumatic arms [28, 29, 30, 31, 20, 32], which offer infinite degrees of freedom (DoFs) and desirable traits such as flexibility and low cost, often using silicone rubber or origami structures. However, their uncontrolled DoFs reduce precision and payload capacity. To overcome this, rigid-soft hybrid arms have been developed, combining rigid bones and joints with soft pneumatic actuators. This reduces uncontrolled compliance by limiting it to the joints. For instance, Ohta et al. created a rigid-bone arm actuated by McKibben actuators [33]. While lightweight and accurate, only five of its seven DoFs are actuated by McKibbens, and the low contraction ratio of McKibben actuators introduces a tradeoff between joint force and range of motion – placing the actuator attachment point closer to the joint would increase range of motion, but decrease the length of the force lever arm, and vice versa. Therefore, McKibben actuator-based arms with limited DoFs find it difficult to achieve both large payload capacity and workspace size. Adding more independently controlled, stronger actuators could help, but will increase design complexity due to the extra tubing and regulators required. Bruder et al. addressed payload limitations with antagonistic stiffening in a McKibben-based soft arm [34], improving force transmission. However, each segment uses only one actuator for stiffening, requiring all DoFs in the segment to stiffen together, which limits motion during stiffening. Origami-based hybrid designs are also promising. Liu et al. used water bomb origami in a rigid-soft arm [35], achieving high payload capacity. Yet, the locking mechanism only engages when fully extended, limiting its practicality. Oh et al. used origami bellows as actuators [36], resulting in strong force output and motion range, but limited DoFs and scaling challenges due to bellow size and pressure constraints. Other approaches focus on high-density actuators. Zhou et al. developed a compact segment with many independently controlled rubber bellows [37], achieving excellent dexterity and precision. However, like all existing designs, it relies on bulky, centralized regulators that reduce portability and hinder use in unstructured environments. In summary, while existing designs address some challenges, none fully overcome all limitations of pneumatic soft arms in unstructured environments. A new design approach is needed.
@@
To address the key limitations of existing pneumatically driven soft arms – namely, limited precision, payload capacity, and reliance on bulky external regulators – we propose a new design approach that integrates pressure regulation directly into the actuators. Building on a recent work [25], where the authors introduced the valve-embedded McKibben actuator (VEMA), we present UMArm, a pneumatically driven rigid-soft hybrid robotic arm designed for high performance in unstructured environments.
@@
UMArm is a highly capable robotic arm with 24 independently controlled actuators, untethered operation capability, and the ability to lift a 3 kg payload while weighing only 1.15 kg. Its design combines embedded actuator pressure regulation, a dense arrangement of independent actuators, tunable stiffness, and rigid structural support within a portable form factor. Specifically, this work introduces three key hardware innovations:
@@
To demonstrate UMArm’s effectiveness, we experimentally validated its ability to perform tasks that have traditionally been challenging for pneumatically driven soft arms. First, a series of experiments is conducted to demonstrate UMArm’s precision, high payload capacity, and untethered capability. The precision of UMArm is showcased through a way point tracking experiment, where it achieves to track given way points with 2.6mm RMSE. Then, its weight-lifting ability is tested with up to 3 kg of payload. Additionally, we tested its untethered operation capability by powering it with a portable compressed air tank, which successfully sustained the UMArm for 12 minutes at full capacity. Second, through the use of antagonistic stiffening, UMArm achieves high-precision manipulation. It can stack three small objects with great consistency. Moreover, the antagonistic stiffness is independently controlled for each joint, granting UMArm the ability to perform directional compliance control. We have shown that the UMArm can structurally stiffen in one direction while remaining soft in another, enabling more adaptive interactions. Third, we demonstrate that UMArm can be worn and used on worksites, highlighting both its portability and its capability of using its compliant structure to compensate for misalignment and disturbances. Finally, to showcase its versatility, we added legs between arm segments, transforming UMArm into an untethered inchworm robot, further broadening its potential application space.
@@ II Design
Figure 2: A: Photo of the robot arm overlaying the annotation of the revolute axes. There are 12 revolute joints in total for all three segments, controlled by 24 actuators. B: Close-up rendered picture of a joint. A U-joint connects two adjacent rigid links. Each U-joint is modeled as 2 revolute joints. C: Rendered Picture showing the structure of one segment of UMArm. There are two groups (of 4) of actuators. Each group controls the two revolute joints within one U-joint. Two groups are stacked reversely with 45° offset angle.
@@
In designing the UMArm, we aim to explore a method for scaling up the number of independently controlled actuators within a compact system and to determine what performance gains can be achieved through this approach. UMArm is a hybrid rigid-soft arm. UMArm has a rigid spine driven by soft actuators. The rigidity of a spine significantly improves precision and force output, and the soft actuators maintain the structural compliance of the arm. As illustrated in figure 1D, the UMArm consists of three structurally identical spine segments. Each segment has three rigid-body links — an upper plate, a central rod, and a lower plate — connected in series by two universal joints (U-joints). We chose to use U-joints because they are lightweight and have no backlash. The U-joints also restrict z-axis rotation, which is helpful since we do not have actuators to control that degree of freedom. Each U-joint (Fig. 2B), can be simplified as 2 revolute joints with perpendicular axes. As a result, the UMArm can be seen as a robot with 12 revolute joints, assuming that the links between the two joints are rigid.
@@
To let the system accommodate a greater number of independently controlled actuators, we employ the valve-embedded McKibben actuator (VEMA) that features self-pressure regulating [25]. All VEMAs are connected to the shared constant pressure source, which can be a single pressurized tube with a larger diameter that runs through the arm. Therefore, we eliminated the need for individual long tubing that connects each actuator to its external regulator. Compared to using the thin, long individual tubing for every actuator, the use of shared, larger-diameter tubing will decrease the whole system’s weight and complexity and improve pressure control performance by lowering the pneumatic friction. Also, after establishing the shared pressure line, adding more actuators to the arm would need very little additional hardware, making it possible to scale up the number of actuators in our arm. Under this pressure bus line architecture, we have installed 24 individually controlled VEMAs to the arm. Having such a large number of independently controlled actuators gives us the ability to simultaneously control the joint’s angle and structural stiffness for all 12 joints. Structural stiffness control is achieved through antagonistically placing the actuators for each joint. In the following sections of this manuscript, we will demonstrate that being able to control the stiffness of every joint can improve payload capacity, arm precision, and safety while working around humans.
@@
Limited joint range of motion remains a significant challenge for McKibben-based antagonistic stiffening arms. McKibben actuators can only contract by approximately 34% [38]. When used to drive a revolute joint on an arm, one can increase the range of motion of the joint by decreasing the length of the lever arm, but at the cost of decreasing torque output. UMArm can address this limitation by incorporating more joints within the same arm length, because although each joint would have a small range of motion (but larger force), the net range of motion across all the joints can remain large. Figure 2C shows the actuator placement. We integrated 8 VEMAs into one arm segment. Four VEMAs control the two joint angles between the upper link and the center link. In the resting state, these actuators are angled toward the center. This ensures better force output at a larger joint angle, where the contracted actuator will have better alignment with the force lever arm, and the elongated actuator will generate larger forces at a given pressure, in accordance with the McKibben model in [39]. The gap between the four upper VEMAs can accommodate another group of four VEMAs that drive the joint angles between the center rod and lower plate. As shown in Fig.2C, the lower group is reversed and rotated by 45° to prevent interference while operating. This configuration of VEMA placement efficiently utilizes the space, thus increasing the density of joints, which, as a result, effectively improves the overall range of motion. The UMArm is built with 3 of these segments. The base segment is designed to withstand more load. It has a longer force lever arm for the VEMAs and uses a variant of VEMA with a larger diameter, which will output larger force compared to the smaller diameter VEMAs at the same length and pressure.

### v2

# UMArm: Untethered, Modular, Portable, Soft Pneumatic Arm

@@ Abstract
Robotic arms in modern industries lack adaptability to unstructured environments. Soft pneumatic arms offer better adaptability and human-robot safety but suffer from limited degrees of freedom, precision, payload capacity, and reliance on bulky external regulators. This work presents ”UMArm,” a novel pneumatic rigid-soft hybrid arm that addresses these limitations through densely integrated, self-regulated McKibben actuators on a lightweight rigid spine. The modified actuators incorporate valves and controllers internally, eliminating individual pressure lines and external regulators, significantly reducing system weight and complexity. UMArm achieves full untethered operation, high payload capacity, precision, and tunable compliance. Portability is demonstrated via wearable assistive arm experiments, and versatility through reconfiguration into an inchworm robot. Results show that high-degree-of-freedom, external-regulator-free pneumatic arm systems like UMArm have great potential for real-world unstructured environments.
@@ I Introduction
Robotic arms are essential components in modern industries. Rigid robotic arms can operate with great precision, speed, and payload capacity, making them ideal for highly structured environments like factories [almurib2012review, brogaardh2007present, kemp2007challenges]. However, their application in unstructured real-world environments remains limited due to their lack of compliance. In unstructured environments, compliance is crucial for safe interaction and effective handling of unexpected disturbances. Compliance can be achieved either by structure or by control. Rigid robots can simulate compliance through control, but it requires the coordinated control of all actuators, and will be limited by controller bandwidth and sensor latency. Without full-body compliance, rigid arms have to operate on highly precise and low-latency spatial information—a requirement that is difficult to satisfy in unstructured environments [zheng2020robust, wang2022control]. Consequently, rigid arms can cause damage to target objects, nearby humans, or the robot itself [yang2019collaborative, zhang2020modular, le2020challenges].
@@
Unlike their rigid counterparts, soft robotic arms have structural compliance. They are capable of mitigating external disturbances and compensating for misalignment through elastic deformation of their structure, making them more suitable for operation in unstructured environments [trimmer2013soft, rus2015design, li2024disturbance]. Their inherent compliance also ensures safety in physical human-robot interaction, enabling use cases such as daily assistance and elder care [ansari2017towards, firth2022anthropomorphic].
@@
A variety of soft actuation strategies have been explored to enable compliance and safety in unstructured environments. HASEL actuators achieve high strain and power output via electrohydraulic zipping, enabling fast, modular actuation; however, their high-voltage requirements and structural fragility pose challenges for unstructured deployment [mitchell2019easy]. Shape memory alloy (SMA)-driven soft actuators offer compact and lightweight actuation with high power-to-weight ratios and low voltage requirements but suffer from slow thermal response, nonlinear behavior, and sensitivity to ambient temperature—factors that hinder precise control [yang2018design]. Hydraulic soft actuators provide high force output[feng2021design], yet their practicality is limited by their bulky fluidic components and slow responses. Tendon-driven designs have also gained attention, offering enhanced workspaces and tunable compliance through mechanisms like trimmed helicoids [guan2023trimmed], but these systems remain vulnerable to friction, slack, and lack robustness under dynamic conditions [guan2023trimmed, yang2018design].
TABLE I: Comparison of UMArm with similar-scale soft robotic arm designs from the literature
Work Untethered DoFs Weight / Force Output Position Control Accuracy
Poly-Limb (Nguyen et al.)[Nguyen2019Soft] Tethered $9_{\text{continuum}}$ 1.6 kg(Arm only) / 0.3 kg payload No closed loop position control
Lightweight PAM Arm (Ohta et al.) [ohta2018design] Tethered $5_{\text{Soft}}+2_{\text{Rigid}}$ 1.2 kg(Arm only) / 0.5 kg payload Joint level control only
Stiffening Arm (Bruder et al.) [bruder2023increasing] Tethered $9_{\text{continuum}}$ 1 kg(Arm only) / 0.5 kg payload No closed loop position control
Koopman Arm (Bruder et al.) [bruder2021koopman] Tethered $6_{\text{Variable Stiffness}}$ 1 kg(Arm only) / 0.275 kg payload Koopman-MPC, 30mm RMSE
UMArm (this work) Untethered $2_{\text{Variable Stiffness}} 1.1$5 kg (with pressure regulators) / 53.3N blocked force Closed loop inverse kinematics control, 0.85 mm RMSE
@@
Among various soft actuation strategies, pneumatic actuators are promising due to their smooth, continuous deformation, high force-to-weight ratio, and compliance in every direction. Unlike cable-driven mechanisms, pneumatic actuators do not require complex cable routing, which simplifies design and thus enhances reliability. Compared to dielectric elastomers, HASEL actuators, and shape memory alloys, pneumatic actuators offer a more practical path toward building human-arm-scale systems that avoids the need for high-voltage operation, complex material packaging, or thermally limited cycle rates. These properties, combined with their resilience to impact, ease of fabrication, and cost-effectiveness, make pneumatic actuators particularly well-suited for building robot arms for deploying into real-world, unstructured environments [whitesides2018soft, xavier2022soft, garcia2024design, galloway2016soft, trivedi2008soft, seleem2023recent].
@@
Although pneumatically actuated soft arms have desirable characteristics, certain drawbacks limit their performance. Firstly, although many pneumatic soft arms are lightweight, they require bulky external pressure regulators [sokolov2023design]. The cumbersome external hardware limits their portability. Secondly, high dexterity pneumatic soft arms need many independently actuated degrees of freedom (DoF), but the centralized, external regulating scheme makes it difficult to increase independent DoFs. This is because adding independently controlled pneumatic actuators requires extra external regulators and more tubing running from the actuators to the external regulators. Large amounts of tubing will introduce extra weight, thereby affecting the payload capacity of the arm. At the same time, long thin tubes increase fluidic resistance, causing pressure to propagate more slowly from the regulator to the actuator, affecting pressure control performance [zuo2024embedded]. Thirdly, structurally soft pneumatic arms might suffer from uncontrolled compliant degrees of freedom (DoF). Under heavy load, they might buckle at unwanted positions, affecting precision and payload capacity [bruder2020data]. Moreover, robots might oscillate at the uncontrolled compliant joints when given aggressive control inputs [li2021position].
Researchers have explored various pneumatic soft arm architectures to address common limitations. Continuum soft pneumatic arms [hongwei2022kinematic, alessi2023ablation, zhu2024design, jiang2020design, garcia2024design, ian2006octarm], which offer infinite degrees of freedom (DoFs) and desirable traits such as flexibility and low cost. They often use silicone rubber or origami structures. However, their uncontrolled DoFs reduce precision and payload capacity.
To overcome this, rigid–soft hybrid arms have been developed, combining rigid bones and joints with soft pneumatic actuators. This reduces uncontrolled compliance by limiting it to the joints. For instance, Ohta et al. created a rigid-bone arm actuated by McKibben actuators [ohta2018design]. While being lightweight, only five of its seven DoFs are actuated by McKibbens, and the low contraction ratio of these actuators introduces a tradeoff between joint force and range of motion. As a result, McKibben-based arms with limited DoFs struggle to achieve both large payload capacity and workspace size. Adding more independently controlled actuators could help, but increases design complexity due to additional tubing and regulators.
@@
Several hybrid designs have also been proposed to improve force transmission and scalability. Bruder et al. addressed payload limitations with antagonistic stiffening in a McKibben-based soft arm [bruder2023increasing], improving force transmission. However, each segment uses only one actuator for stiffening, requiring all DoFs in the segment to stiffen together, which limits motion during stiffening. Origami-based hybrid designs are also promising. Liu et al. used water bomb origami in a rigid–soft arm [liu2023design], achieving high payload capacity, though the locking mechanism only engages when fully extended. Oh et al. used origami bellows as actuators [oh2024hybrid], producing strong force output and motion range, but with limited DoFs and scaling challenges. Zhou et al. developed a compact segment with many independently controlled rubber bellows [zhou2024pneumatic], achieving excellent dexterity and precision. However, the scale of the robot is relatively small, the possibility of scaling up is unknown, and, like all existing designs, it relies on bulky, centralized regulators that reduce portability and hinder use in unstructured environments.
@@
To address the key limitations of existing pneumatically driven soft arms – namely, limited precision, payload capacity, and reliance on bulky external regulators – we propose a new design approach that integrates pressure regulation directly into the actuators. Building on a recent work [zuo2024embedded], where the authors introduced the valve-embedded McKibben actuator (VEMA), we present UMArm, a pneumatically driven rigid-soft hybrid robotic arm designed for high performance in unstructured environments.
@@
UMArm is a highly capable robotic arm with 24 independently controlled actuators, untethered operation capability, and the ability to directionally generate a 53.3N force while weighing only 1.15 kg. Its design combines embedded actuator pressure regulation, a dense arrangement of independent actuators, tunable stiffness, and rigid structural support within a portable form factor. Specifically, this work introduces three key hardware innovations:
@@
To better situate our contribution, Table I summarizes representative soft robotic arms from the literature, focusing on systems of comparable scale to UMArm. Early continuum and McKibben-based designs demonstrated impressive compliance, dexterity, or payload capacity, but lacked closed-loop precision and remained tethered to external regulators. More recent works introduced stiffening mechanisms or advanced control strategies (e.g., Koopman-based MPC), which improved load handling and tracking but still relied on bulky hardware and achieved only centimeter-level endpoint accuracy. In contrast, UMArm is fully untethered, portable, and precise (0.85 mm RMSE), with high force output and 12 actively controlled DoFs with tunable stiffness. This combination of features distinguishes UMArm from prior soft robotic arms and enables high-performance operation in unstructured environments.
To demonstrate UMArm’s effectiveness, we experimentally validated its ability to perform tasks that have traditionally been challenging for pneumatically driven soft arms. First, we conducted a series of experiments to evaluate UMArm’s precision, force-generation capability, and untethered operation. Precision was assessed through a position-control step response experiment and a trajectory-tracking experiment. The step response experiment achieved a positional control RMSE of 0.85 mm, while the trajectory-tracking experiment showed a near-identical match between the target and actual trajectories. UMArm’s force-generation ability was demonstrated with a tested output up to 53.3 N. Additionally, we tested its untethered operation capability by powering it with a portable compressed air tank, which successfully sustained the UMArm for 12 minutes at full capacity. Second, through the use of antagonistic stiffening, UMArm achieves high-precision manipulation. It can stack three small objects with great consistency. Moreover, the antagonistic stiffness is independently controlled for each joint, granting UMArm the ability to perform directional compliance control. We have shown that the UMArm can structurally stiffen in one direction while remaining soft in another, enabling more adaptive interactions. Third, we demonstrate that UMArm can be worn and used on worksites, highlighting both its portability and its capability of using its compliant structure to compensate for misalignment and disturbances. Finally, to showcase its versatility, we added legs between arm segments, transforming UMArm into an untethered inchworm robot, further broadening its potential application space.
@@ II Design
Figure 2: A: Photo of the robot arm overlaying the annotation of the revolute axes. There are 12 revolute joints in total for all three segments, controlled by 24 actuators. B: Close-up rendered picture of a joint. A U-joint connects two adjacent rigid links. Each U-joint is modeled as 2 revolute joints. C: Rendered Picture showing the structure of one segment of UMArm. There are two groups (of 4) of actuators. Each group controls the two revolute joints within one U-joint. Two groups are stacked reversely with $\pi/4$ rad offset angle.
@@
In designing the UMArm, we aim to explore a method for scaling up the number of independently controlled actuators within a compact system and to determine what performance gains can be achieved through this approach. UMArm is a hybrid rigid-soft arm. UMArm has a rigid spine driven by soft actuators. The rigidity of a spine significantly improves precision and force output, and the soft actuators maintain the structural compliance of the arm. As illustrated in figure 1D, the UMArm consists of three structurally similar spine segments, with decreasing length from base to tip. Each segment has three rigid-body links — an upper plate, a central rod, and a lower plate — connected in series by two universal joints (U-joints). We chose to use U-joints because they are lightweight and have no backlash. The U-joints also restrict z-axis rotation, which is helpful since we do not have actuators to control that degree of freedom. Each U-joint (Fig. 2B), can be simplified as 2 revolute joints with perpendicular axes. As a result, the UMArm can be seen as a robot with 12 revolute joints, assuming that the links between the two joints are rigid.
@@
To let the system accommodate a greater number of independently controlled actuators, we employ the valve-embedded McKibben actuator (VEMA) that features self-pressure regulating [zuo2024embedded]. All VEMAs are connected to the shared constant pressure source, which can be a single pressurized tube with a larger diameter that runs through the arm. Therefore, we eliminated the need for individual long tubing that connects each actuator to its external regulator. Compared to using the thin, long individual tubing for every actuator, the use of shared, larger-diameter tubing will decrease the whole system’s weight and complexity and improve pressure control performance by lowering the pneumatic friction. Also, after establishing the shared pressure line, adding more actuators to the arm would need very little additional hardware, making it possible to scale up the number of actuators in our arm. Under this pressure bus line architecture, we have installed 24 individually controlled VEMAs to the arm. Having such a large number of independently controlled actuators gives us the ability to simultaneously control the joint’s angle and structural stiffness for all 12 joints. Structural stiffness control is achieved through antagonistically placing the actuators for each joint. In the following sections of this manuscript, we will demonstrate that being able to control the stiffness of every joint can improve payload capacity, arm precision, and safety while working around humans.
@@
Limited joint range of motion remains a significant challenge for McKibben-based antagonistic stiffening arms. McKibben actuators can only contract by approximately 34% [bruder2021chain]. When used to drive a revolute joint on an arm, one can increase the range of motion of the joint by decreasing the length of the lever arm, but at the cost of decreasing torque output. UMArm can address this limitation by incorporating more joints within the same arm length, because although each joint would have a small range of motion (but larger force), the net range of motion across all the joints can remain large. Figure 2C shows the actuator placement. We integrated 8 VEMAs into one arm segment. Four VEMAs control the two joint angles between the upper link and the center link. In the resting state, these actuators are angled toward the center. This ensures better force output at a larger joint angle, where the contracted actuator will have better alignment with the force lever arm, and the elongated actuator will generate larger forces at a given pressure, in accordance with the McKibben model in [bruder2018force]. The gap between the four upper VEMAs can accommodate another group of four VEMAs that drive the joint angles between the center rod and lower plate. As shown in Fig.2C, the lower group is reversed and rotated by $\pi/4$ rad to prevent interference while operating. This configuration of VEMA placement efficiently utilizes the space, thus increasing the density of joints, which, as a result, effectively improves the overall range of motion. The UMArm is built with 3 of these segments. The base segment is designed to withstand more load. It has a longer force lever arm for the VEMAs and uses a variant of VEMA with a larger diameter, which will output larger force compared to the smaller diameter VEMAs at the same length and pressure.